what mass of copper (ii) oxide is required to produce 12.8 g of copper?
1
Expert's answer
2019-06-07T12:56:45-0400

"m(CuO)=\dfrac{m(Cu)*M(CuO)}{M(Cu)}=\dfrac{12.8g*79.55g/mol}{63.55g/mol}=16.02g"